OpenStudy (anonymous):

Given the parent functions f(x) = log10 x and g(x) = 3x - 1, what is f(x) g(x)? (2 points) f(x) g(x) = log10 (3x - 1)x f(x) g(x) = log10 x3x - 1 f(x) g(x) = 3x log10 x + log10 x f(x) g(x) = log10 x - 3x log10 x

OpenStudy (jiteshmeghwal9):

is\[f(x)=\log_{10}x\] ?

OpenStudy (anonymous):

yes

OpenStudy (jiteshmeghwal9):

\[f(x).g(x)=\log_{10}x \times (3x-1)\]\[f(x).g(x)=3xlog_{10}x -\log_{10}x\]

OpenStudy (anonymous):

thanks so much!!<3

OpenStudy (jiteshmeghwal9):

Is 2nd option correct?

OpenStudy (usukidoll):

maybe it's just \[f(x).g(x)=\log_{10}x \times (3x-1) \] ????

OpenStudy (jiteshmeghwal9):

yeah I think so .

OpenStudy (usukidoll):

strange because you can just leave it as that or distribute.